OpenStudy (anonymous):

equation of tangent line y=1-x/1+x at x=0

OpenStudy (callisto):

Differentiate the equation with respect to x, then evaluate the derivative at x=0 to get the slope of the tangent line. Get the y-coordinates at x=0, i.e. sub. x=0 into the equation. You have a point, and the slope of the tangent line, you get find the equation.

OpenStudy (anonymous):

\[y=\frac{ 1-x }{ 1+x }\] ??

OpenStudy (anonymous):

\[\frac{ d }{ dx }(\frac{ 1-x }{ 1+x })= \frac{ -2 }{ (x+1)^2 }\]

OpenStudy (anonymous):

now substitute for x=0

OpenStudy (anonymous):

so \[y=-2\]

OpenStudy (amistre64):

btw, y does not equal -2